# ProFL: Performative Robust Optimal Federated Learning

### v1

@@ 6 Experiments
Figure 1: Effectiveness of Proposed methods
@@
Figure 2: Dynamic contribution with Different Sample Sizes, Learning Rates, and $H$
Comparison to Centralized Setting. Table 1 shows that adapting performative gradients to the FL framework improves performance. There are 10 clients with $\gamma_{i}\in[\overline{\gamma}-\alpha,\overline{\gamma}+\alpha]$, where larger $\alpha$ indicates more heterogeneity. As client heterogeneity increases, the performance of both algorithms degrades. Under homogeneous distribution shifts ($\alpha=0$), both algorithms perform similarly. ProFL (ours) outperforms PG when $\alpha=0.25$ or $\alpha=0.5$, as the centralized approach struggles to handle variance shifts by estimating a single $\frac{df}{d\theta}$ on the server.
Effectiveness of our proposed methods. Fig. 1(a) shows that contamination can significantly degrade performance, but the proposed robust gradient can effectively reduce the impact of contaminated data when computing the gradients i.e., the performance with the robust gradient in our method is almost the same as the case when outliers do not exist at all. This case study, pricing with dynamic demands, involves 10 heterogeneous clients with different $D_{i}(\theta)$, each encounters an $\epsilon_{i}$ fraction of contaminated data from a fixed but unknown exogenous distribution $Q_{i}$. Although we present the results when $Q_{i}\sim\mathcal{N}(\mu_{o},\sigma^{2}_{o})$, we observed similar results when $Q_{i}$ follows other distributions and are different among clients.
Fig. 1(b) shows that with a limited sample size and many clients ($n_{i}=60,|\mathcal{V}|=100$), estimating $\widehat{\frac{df_{i}}{d\theta}}$ on the server side results in a lower loss value compared to local estimation and approaches the optimal value. Fig. 1(c) demonstrates the effectiveness of adaptive sample sizing. We tested two tolerances (0.05 for purple, 0.1 for red), with each client processing up to 1000 samples per iteration. The adaptive approach significantly reduces the sample size while achieving similar results within the error tolerance, leading to substantial computational savings. Both experiments are based on the case study pricing with dynamic contribution, involving a non-linear $f_{i}(\theta)$.
@@
Table 2: Accuracy of Binary Classification on Synthetic and Realistic Datasets
DATASET POFL PROFL
Same Distribution 88.00 $\pm$ 2.69 99.23 $\pm$ 0.29
Different Distributions 62.44 $\pm$ 0.52 92.50 $\pm$ 0.29
Credit Dataset 94.73 $\pm$ 0.00 94.73 $\pm$ 0.00
Adult Dataset 54.74 $\pm$ 0.02 60.78 $\pm$ 0.81
Convergence to the performative optimal solution. Table 2 shows that ProFL outperforms PFL in the case study of binary strategic classification with 10 heterogeneous clients. Since PFL cannot handle contaminated data, we set $\epsilon_{i}=0$ for these experiments. The first two rows show that ProFL is more stable and significantly more accurate than PFL on Gaussian synthetic data. PFL converges to a PS point rather than a PO point. On the Give Me Some Credit dataset (third row), both algorithms converge to the same point, but ProFL has a faster convergence rate (see Fig. 7(c) in the Appendix). With the Adult dataset, ProFL achieves $60.78\%$ accuracy, outperforming PFL’s $54.74\%$. These results highlight ProFL’s practical advantages in both convergence speed and accuracy. More results are in Appendix D.1.
Appendix A Algorithms and Details
A.1 Pseudocode of the PoFL
Algorithm 2 presents the pseudocode of the PoFL algorithm discussed in the main paper.
Algorithm 2 PoFL
1:  Input: Learning rate $\eta$, estimation window $H$, function $\widehat{f}$, total iterations $T$, local update iterations $R$.
2:  for $t=0$ to $T-1$ do
4:     for $i\in\mathcal{I}_{t}$ in parallel do
6:        if $t\leq H$ then
7:           Compute estimator $\widehat{f}\left(\mathcal{S}_{i}\right)$ of $f^{t}_{i}$.
8:           Update $\theta^{t+1}_{i}\leftarrow\text{Proj}_{\Theta}\left(\theta^{t}_{i}-\eta\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i})\right)$.
9:        else
10:           Compute estimator $\widehat{f}\left(\mathcal{S}_{i}\right)$ of $f^{t}_{i}$.
11:           Form matrices $\Delta\theta_{i}$ and $\Delta f_{i}$ and compute $\Delta f_{i}(\Delta\theta_{i})^{\dagger}$.
12:           $\theta^{t+1}_{i}\leftarrow\text{Proj}_{\Theta}\left(\theta^{t}_{i}-\eta\left(\widehat{\nabla}\mathcal{L}_{i,1}(\theta^{t}_{i})+\widehat{\nabla}\mathcal{L}_{i,2}(\theta^{t}_{i})\right)\right)$.
13:     If $t\mod R=0$, client $i\in\mathcal{I}_{t}$ sends $\theta^{t}_{i}$ to the server and server updates the model $\overline{\theta}^{t+1}=\sum_{i\in\mathcal{I}_{t}}\alpha_{i}\theta^{t+1}_{i}$.
14:  Output: $\overline{\theta}^{T}$
Algorithm 3 Robust Gradient
1:  Input: Set $\mathcal{S}_{i}$ of samples $(z_{j})_{j=1}^{n_{i}}$, thresholds $C,J\in(0,1)$, model $\theta^{t}_{i}\in\Theta$
2:  Let ^$\widehat{\nabla}\leftarrow\frac{1}{|\mathcal{S}_{i}|}\sum_{z_{j}\in\mathcal{S}_{i}}\nabla\ell(z_{j};\theta^{t}_{i})$.
3:  Let $\left[\nabla\ell(z_{j};\theta^{t}_{i})-\widehat{\nabla}\right]_{z_{j}\in\mathcal{S}_{i}}$ be the $|\mathcal{S}_{i}|\times d$ matrix of centered gradients.
4:  Apply SVD to this matrix and find top right singular vector $v$.
5:  Compute the outlier score $\tau_{j}$ of $\nabla\ell(z_{j};\theta^{t}_{i})$ by Eq. (9).
6:  Divide the interval $[0,\max\tau_{j}]$ into $B$ equal-length segments. Let $\phi_{k}$ be the number of $\tau_{j}$ in the $k$-th segment.
7:  Find the smallest $k\in\{1,\cdots,B\}$ with $\phi_{k}<C\cdot|\mathcal{S}_{i}|$. Set $\phi$ as the lower bound of the $k$-th segment.
9:  Compare the average gradients $\widehat{\nabla}$ of $\mathcal{S}_{i}$ and $\widehat{\nabla}^{\prime}$ of $\mathcal{S}^{\prime}_{i}$. If ^^^$\frac{\|\widehat{\nabla}-\widehat{\nabla}^{\prime}\|}{\|\widehat{\nabla}\|}<J$ or $|\mathcal{S}^{\prime}_{i}|\leq\frac{n_{i}}{2}$, return $\mathcal{S}^{\prime}_{i}$ and $\widehat{\nabla}^{\prime}$; otherwise, set $\mathcal{S}_{i}\leftarrow\mathcal{S}^{\prime}_{i}$ and repeat from line 2.
A.2 Outlier removal mechanism for robust gradient
To estimate $\nabla\mathcal{L}_{i,1}$, we should average $\nabla\ell(z_{j};\theta_{i}^{t})$ over samples drawn from $D_{i}({\theta}_{i}^{t})$. Because the samples collected by client $i$ in round $t$ are drawn from $P_{i}(\theta_{i}^{t})$, they are noisy, with an unknown $\epsilon_{i}$ fraction of outliers from an unknown distribution $Q_{i}$. ProFL needs to identify contaminated data and eliminate the corresponding contaminated gradients. We propose a mechanism that can remove these contaminated gradients and estimate gradients that are robust to outliers. Importantly, ProFL is independent of the specific outlier identification and removal mechanism used and can function with any such mechanism. In our paper, we assume that contaminated data can follow an arbitrary distribution. In practice, additional information about the contamination process may be available based on the application or context, allowing us to design more efficient (and less expensive) outlier identification mechanisms leveraging this knowledge.
As shown in Algorithm 3, our mechanism takes noisy samples $\mathcal{S}_{i}$ and current local model $\theta_{i}^{t}$ as inputs. The mechanism first computes the gradients $\nabla\ell(z_{j};\theta_{i}^{t})$ of all samples and iteratively identifies and removes the contaminated gradients.
Since both the distribution $Q_{i}$ and fraction $\epsilon_{i}$ of contaminated data are unknown, we consider gradients contaminated if they exhibit two crucial properties: 1) they have large effects on the learned model and can disturb the training process (i.e., gradients have large magnitudes and systematically point in a specific direction); 2) they are located farther away from the average in the vector space (i.e., “long tail” data). To detect gradients that satisfy these properties, we adapt the approach in Shan et al., (2023), which leverages singular value decomposition (SVD). Instead of preselecting the threshold for the sum of all outlier scores $\tau_{j}$, which is difficult to determine without any prior knowledge of the data, our threshold is based on the two crucial properties, making it much easier to predict the values of $C$ and $J$. Specifically, we construct a matrix $[\nabla\ell(z_{j};\theta^{t}_{i})-\widehat{\nabla}]_{z_{j}\in\mathcal{S}_{i}}$ using the centered gradients and find the top right singular vector $v$, where $\widehat{\nabla}$ represents the average of all gradients (lines 2-4). The centered gradients that are closer to $v$ are more likely to be outliers and we assign each gradient an outlier score $\tau_{j}$ as follows:
Given the outlier scores $\{\tau_{j}\}$, the gradients with scores above a threshold $\phi$ are considered contaminated and are discarded. Unlike Shan et al., (2023) that uses a fixed pre-defined threshold, our method finds $\phi$ automatically based on the outlier score distribution. Since we assume the outliers are in the “long tail” of the distribution, we divide the interval $[0,\max\tau_{j}]$ into $B$ equal-length segments and compute the relative frequency distribution of scores. $\phi$ is set as the smallest segment with the relative frequency below a threshold (lines 6-7). For the remaining gradients with $\tau_{j}<\phi$, we compute their average $\widehat{\nabla}^{\prime}$ (lines 8-9).
Appendix B Additional Related Work
B.1 Performative Prediction in Centralized Settings
Performative prediction (Perdomo et al.,, 2020) was first introduced in 2020 corresponding to the optimization framework to deal with endogenous data distribution shifts where the model deployed to the environment will affect the subsequent data, resulting in the collected data changing as the deployed model changes. Common applications of performative prediction include strategic classification(Hardt et al.,, 2016; Xie and Zhang,, 2024) and predictive policing (Ensign et al.,, 2018). Perdomo et al., (2020) first proposes an iterative optimization procedure named Repeated Risk Minimizationto find a performative stable point and also bounded the distance between the performative stable point and the performative optimal point. Mendler-Dünner et al., (2020) designed the first algorithm to find the performative stable point under the online learning setting. Miller et al., (2021) first solved the problem of performative prediction by directly optimizing performance risk to find the performative optimal point, but the scope of the distribution map is restricted. Mendler-Dünner et al., (2022) proved the convergence of greedy deployment and lazy deployment after each random update under the assumption of smoothness and strong convexity and gave the convergence rate. Izzo et al., (2021) also designed an algorithm performative gradient descent for finding the performative optimal solution. Zhao, (2022) relaxes the assumption of strong convexity to the weakly-strongly convex case and proves the convergence of the deployment. Somerstep et al., (2024); Bracale et al., (2024) focused on learning the distribution map in performative prediction and proposed a framework named Reverse Causal Performative Prediction. In addition, performative prediction is also related to reinforcement learning(Zheng et al.,, 2022) and bandit problems(Chen et al.,, 2024).
B.2 Performative Prediction in Distributed System
To the best of our knowledge, Jin et al., (2024) is the only work discussing performativity under the distributed setting, where they generalized Fedavg to P-Fedavg and proved the uniqueness of the performative stable solution found by the algorithm. They also quantified the distance between the performative stable solution and the performative optimal solution. While the work is a breakthrough as the first to take model-dependent distribution shifts into account under the FL setting, it is only a strict generalization of Perdomo et al., (2020) without the ability to find the performative optimal point and deal with the contaminated data.
B.3 Federated Learning under Dynamic Data
Previous works (Tan et al.,, 2023; Li et al., 2020a, ) pointed out that the training paradigm of FedAvg may violate the i.i.d. assumptions and cause some feature classes to be over-represented while others to be under-represented. Li et al., 2020b ; Li et al., (2021) steer the local models towards a global model by adding a regularization term to guarantee convergence when the data distributions among different clients are non-IID. Li et al., 2020b proposes FedProx as a solution. Li et al., (2021) aims at obtaining good local models on clients rather than a global consensus model. Meanwhile, another line of work focused on dealing with the statistical heterogeneity by clustering (Briggs et al.,, 2020; Ghosh et al.,, 2020; Sattler et al.,, 2020) by aggregating clients with similar distribution into the same cluster. While most previous works assumed the “dynamic” is among different clients, another line of literature focuses on FL in a dynamic environment where various distribution shifts occur. Some works aim to deal with time-varying contribution rates of clients with local heterogeneity (Sattler et al.,, 2021; Park et al.,, 2021; Wang and Ji,, 2022; Zhu et al.,, 2021).
B.4 Huber’s $\epsilon$-contamination Model
Proposed by Huber, (1992), Huber’s epsilon contamination model serves as a framework for robust statistical analysis with outliers and contaminations in the dataset. This model assumes that besides the data distribution we aim to learn, a small proportion of samples can come from an arbitrary distribution. This contamination framework is crucial for creating machine learning algorithms robust to outliers, thereby ensuring more reliable analysis in practical scenarios where data imperfections are common. Huber’s model has been fundamental in the field of robust statistics, influencing a wide range of applications and subsequent research. Specifically, denote $D$ as the true data distribution and $Q$ as a random distribution. $P=(1-\epsilon)D+Q$ is the distribution with some corruption $Q$ and the portion of corruption data is $\epsilon$. The local dataset of client $i$ is $P_{i}=(1-\epsilon_{i})D_{i}(\theta_{i})+\epsilon_{i}Q_{i}$.
Appendix C Additional Proofs not Detailed in the Main Paper
C.1 Proof of Theorem 5.10
@@ Appendix C Additional Proofs not Detailed in the Main Paper
The first inequality arises from the application of Jensen’s inequality to two terms. The second inequality arises from Assumption 5.2. The last inequality arises form Lemma C.1.
@@
Denote the empirical result of the performative gradient of client $i$ in global iteration $t$ and local iteration $r$ as ^$\hat{\nabla}\mathcal{L}_{i}(\theta_{i}^{t})=\hat{\nabla}\mathcal{L}_{i,1}(\theta_{i}^{t})+\hat{\nabla}\mathcal{L}_{i,2}(\theta_{i}^{t})$. According to Lemma C.2, we have
@@
where $\text{err}^{t}_{i}$ is the error result in sampling of client $i$ on iteration $t$ and $[\text{err}^{t}_{i}]=0$ for all $t\in[0,T-1].$ Similarly we have denote $\theta^{t}_{i}$ as $\theta_{t}$ , $\theta^{t-1}_{i}$ as $\theta_{t-1}$ , $\text{err}^{t}_{i}$ as $\text{err}_{t}$ , and $\text{err}^{t-1}_{i}$ as $\text{err}_{t-1}.$
@@
Where .$\left\|.\right\|_{F}$ denotes the Frobenius norm. (12) and (15) arise from the property that $\mathbb{E}\left[\text{err}_{t}\right]=0$ for all $t\in[0,T]$ and $\text{err}_{t}$ is independent to $a_{k}$ and $\frac{df}{d\theta}$. (13) arises from Jensen’s inequality of two terms. (14) arises form the properties of matrix norm that if $A$ and $B$ are two matrix $\left\|AB\right\|\leq\left\|A\right\|\left\|B\right\|$ and $\left\|A\right\|\leq\left\|A\right\|_{F}.$ (16) arises form $\left\|\Delta\theta^{\dagger}\right\|\leq\left\|\Delta\theta\right\|^{-1}\leq\lambda^{-1}_{i,min}$, where $\lambda_{i,min}$ is the smallest singular values of $\Delta\theta$ of client $i$ for all $t\in[0,T]$ and $\left\|\text{err}\right\|^{2}_{i}$ is the maximum of $\left\|\text{err}_{t}\right\|^{2}$ for all $t\in[0,T]$. Because $\left\|\text{err}_{t}\right\|^{2}$ depends on the data distribution, the estimation methods, the number of samples $n_{i}$, as well as the probability $\varphi_{i}$ of $\mathbf{Pr}(\left\|\text{err}_{t}\right\|\leq 1-\varphi_{i})$, we use $f_{\varphi}(\varphi_{i};n_{i})$ to represent this term.
@@
The second part will decrease to zero as $n_{i}\rightarrow 0.$ The upper bound of this part is $\mathcal{O}\left(\ell^{2}_{\max}dF^{2}f_{\varphi}(\varphi_{i};n_{i})\right)$ by using Lemma 12 in Izzo et al., (2021). Then we can obtain
@@
(19) arises from Eq. (3). (20) arises from Lemma 10 of Izzo et al., (2021) and Lemma 5.6. (21) arises from Jensen’s inequality of two terms. (22) arises from Lemma 5.7. ∎
C.4 Additional Lemmas and their Proofs
Lemma C.1.

### v2

@@ 5.2 Numerical Results
Table 1: Performative risk of ProFL (federated) and PG (centralized) under varying client heterogeneity ($\alpha$). ProFL outperforms PG as heterogeneity increases.
@@
Figure 2: (a) Robust gradient mitigates the impact of data contamination. (b) Estimation of $\frac{df_{i}}{d\theta}$ on the server outperforms local estimation in the large-client, limited-sample setting ($|\mathcal{V}|=100$, $n_{i}=60$). (c) Adaptive sampling (tolerances 0.05/0.1) reduces sample usage while maintaining accuracy.
@@
Figure 3: (a) Increasing sample size $n_{i}$ improves stability and convergence to the PO point. (b) A smaller learning rate $\eta$ slows convergence but yields solutions closer to the PO point. (c) A smaller estimation window $H$ produces more local estimates of $\frac{df_{i}}{d\theta}$, reducing final error at the cost of slower convergence.
Effectiveness of our proposed methods. Fig. 2(a) shows that contamination can significantly degrade performance, but the proposed robust gradient can effectively reduce the impact of contaminated data when computing the gradients i.e., the performance with the robust gradient in our method is almost the same as the case when outliers do not exist at all. Each client encounters an $\epsilon_{i}$ fraction of contaminated data from a fixed but unknown exogenous distribution $Q_{i}$. We select $B=50$, $J_{1}=0.01$, and $J_{2}=0.05$ for the robust gradient method. Although we present the results when $Q_{i}\sim\mathcal{N}(\mu_{o},\sigma^{2}_{o})$, we observed similar results when $Q_{i}$ follows other distributions and are different among clients. Fig. 2(b) shows that with a limited sample size and many clients ($n_{i}=60,|\mathcal{V}|=100$), estimating $\widehat{\frac{df_{i}}{d\theta}}$ on the server side results in a lower loss value compared to local estimation and approaches the optimal value. Fig. 2(c) demonstrates the effectiveness of adaptive sample sizing. We tested two tolerances (0.05 for purple, 0.1 for red), with each client processing up to 1000 samples per iteration. The adaptive approach significantly reduces the sample size while achieving similar results within the error tolerance, leading to substantial computational savings. Both experiments have non-linear $f_{i}(\theta)$.
@@
(a) Same Distributions
(b) Different Distributions
(c) Give Me Some Credit
(d) Adult
Figure 4: Convergence comparison between ProFL and PFL on both synthetic and real-world datasets. ProFL consistently achieves higher accuracy with fewer communication rounds.
Convergence to the performative optimal solution. PFL converges to the PS point $\theta^{\text{PS}}$, which can be suboptimal, whereas ProFL consistently converges closer to the PO point $\theta^{\text{PO}}$. We conduct several experiments across diverse case studies to support this claim. On Gaussian synthetic data, ProFL demonstrates significantly better accuracy and stability than PFL (Figs. 4(a)4(b)), which converges to $\theta^{\text{PS}}$. On real datasets (Figs. 4(c)4(d)), ProFL achieves faster convergence and higher final accuracy (e.g., $60.78\%$ vs. $54.74\%$ on Adult). Since PFL is sensitive to outliers, we set $\epsilon_{i}=0$ for fair comparison. In the non-linear $f_{i}(\cdot)$ cases (Fig. 5(a) and 5(b)), ProFL again converges faster and reaches a point much closer to $\theta^{\text{PO}}$, while PFL remains at the suboptimal $\theta^{\text{PS}}$. Additionally, in Fig. 5(b), ProFL demonstrates more stable convergence despite a persistent gap to $\theta^{\text{PO}}$ caused by the non-linearity of $f_{i}(\cdot)$. Strategies to mitigate this gap are discussed in Subsection 4.3.
(a) Pricing
(b) Regression
Figure 5: Convergence for pricing with dynamic contribution and regression with dynamic contribution. ProFL approaches the performative optimum more closely than PFL, with faster convergence and reduced oscillation, despite a gap due to non-linear $f_{i}(\cdot)$.
(a) Different Heterogeneity
(b) Sample Size
Figure 6: Impact of (a) heterogeneity and (b) sample size on convergence. Only ProFL reaches the PO point.
In Fig. 6 only ProFL reaches the optimal solution. Across all settings, PFL either stagnates at $\theta^{\text{PS}}$ or exhibits oscillations under complex dynamics. Fig. 6(a) illustrates that ProFL maintains convergence to $\theta^{\text{PO}}$ across increasing heterogeneity ($\alpha\in\{0.1,\,0.3,\,1\}$), and Fig. 6(b) shows that increasing sample size improves the stability of both algorithms, although only ProFL ultimately reaches the optimal solution.
6 Conclusion and Future Work
In this work, we proposed ProFL, the first approach to find the PO point in a FL setting. We also analyzed the impact of data contamination, addressing key challenges commonly observed in FL: limited local data sample sizes, constrained local computation, and the potentially high communication cost associated with multiple rounds of model exchange. We provided a convergence analysis under the PL condition, which is weaker than the assumptions required by PFL. However, several limitations remain. Our experiments rely on synthetic distribution shifts due to the lack of real-world datasets with dynamic performative feedback. Future work includes exploring real datasets for more realistic evaluation and relaxing the PL condition to cover broader models such as neural networks, especially large language models.
Appendix A Additional Proofs not Detailed in the Main Paper
A.1 Proof of Theorem 4.10
@@ Appendix A Additional Proofs not Detailed in the Main Paper
The first inequality arises from the application of Jensen’s inequality to two terms. The second inequality arises from Assumption 4.2. The last inequality arises form Lemma A.1.
@@
Denote the empirical result of the performative gradient of client $i$ in global iteration $t$ and local iteration $r$ as $\hat{\nabla}\mathcal{L}_{i}(\theta_{i}^{t})=\hat{\nabla}\mathcal{L}_{i,1}(\theta_{i}^{t})+\hat{\nabla}\mathcal{L}_{i,2}(\theta_{i}^{t})$. According to Lemma A.2, we have
@@
where $\text{err}^{t}_{i}$ is the error result in sampling of client $i$ on iteration $t$ and $\text{err}^{t}_{i}]=0$ for all $t\in[0,T-1].$ Similarly we have denote $\theta^{t}_{i}$ as $\theta_{t}$ , $\theta^{t-1}_{i}$ as $\theta_{t-1}$ , $\text{err}^{t}_{i}$ as $\text{err}_{t}$ , and $\text{err}^{t-1}_{i}$ as $\text{err}_{t-1}.$
@@
Where $\left\|.\right\|_{F}$ denotes the Frobenius norm. (12) and (15) arise from the property that $\mathbb{E}\left[\text{err}_{t}\right]=0$ for all $t\in[0,T]$ and $\text{err}_{t}$ is independent to $a_{k}$ and $\frac{df}{d\theta}$. (13) arises from Jensen’s inequality of two terms. (14) arises form the properties of matrix norm that if $A$ and $B$ are two matrix $\left\|AB\right\|\leq\left\|A\right\|\left\|B\right\|$ and $\left\|A\right\|\leq\left\|A\right\|_{F}.$ (16) arises form $\left\|\Delta\theta^{\dagger}\right\|\leq\left\|\Delta\theta\right\|^{-1}\leq\lambda^{-1}_{i,min}$, where $\lambda_{i,min}$ is the smallest singular values of $\Delta\theta$ of client $i$ for all $t\in[0,T]$ and $\left\|\text{err}\right\|^{2}_{i}$ is the maximum of $\left\|\text{err}_{t}\right\|^{2}$ for all $t\in[0,T]$. Because $\left\|\text{err}_{t}\right\|^{2}$ depends on the data distribution, the estimation methods, the number of samples $n_{i}$, as well as the probability $\varphi_{i}$ of $\mathbf{Pr}(\left\|\text{err}_{t}\right\|\leq 1-\varphi_{i})$, we use $f_{\varphi}(\varphi_{i};n_{i})$ to represent this term.
@@
The second part will decrease to zero as $n_{i}\rightarrow 0.$ The upper bound of this part is $\mathcal{O}\left(\ell^{2}_{\max}dF^{2}f_{\varphi}(\varphi_{i};n_{i})\right)$ by using Lemma 12 in pmlr-v139-izzo21a. Then we can obtain
@@
(19) arises from Eq. (4). (20) arises from Lemma 10 of pmlr-v139-izzo21a and Lemma 4.6. (21) arises from Jensen’s inequality of two terms. (22) arises from Lemma 4.7. ∎
A.4 Additional Lemmas and their Proofs
Lemma A.1.